OpenStudy (anonymous):

describe how you would solve -8 = 1/9 ( -9t+27 )

OpenStudy (hba):

Multiply both sides by 9

OpenStudy (anonymous):

-8 = (1/9)(-9T+27) Divide the right by (1/9) which becomes on the left -8*9=-9T+27 Subtract from the right 27, which become on the left -72-27=-9T Divide both sides by -9 -99/-9=T T= 11